# High-Precision Transition Energy Measurements of Neonlike Fe XVII Ions

### v2

@@ Abstract
We improve by a factor of 4-20 the energy accuracy of the strongest soft X-ray transitions of Fe XVII ions by resonantly exciting them in an electron beam ion trap with a monochromatic beam at the P04 beamline of the PETRA III synchrotron facility. By simultaneously tracking instantaneous photon-energy fluctuations with a high-resolution photoelectron spectrometer, we minimize systematic uncertainties down to 10-15 meV, or velocity-equivalent 5 km/s in their rest energies, substantially improving our knowledge of this key astrophysical ion. Our large-scale configuration-interaction computations include more than four million relativistic configurations and agree with the experiment at a level without precedent for a ten-electron system. Thereby, theoretical uncertainties for interelectronic correlations become far smaller than those of quantum electrodynamics (QED) corrections. The present QED benchmark strengthens our trust in future calculations of many other complex atomic ions of interest to astrophysics, plasma physics, and for the development of optical clocks with highly charged ions.
@@ 1 Introduction
Over the past three decades, extensive research has focused on the soft X-ray emission from Ne-like iron (Fe XVII, Fe16+), particularly in hot astrophysical plasmas observed by Chandra and XMM-Newton (Behar et al., 2001; Brinkman et al., 2001). The dominant spectral transitions and of Fe XVII within the 700 to 850 eV range (14.5 to 17.5 Å) play a crucial role in deducing the plasma parameters across various sources. These parameter include electron temperature, density, elemental abundance, gas motion, and photon scattering opacity (Parkinson, 1973; Smith et al., 1985; Schmelz et al., 1992; Waljeski et al., 1994; Phillips et al., 1996; Behar et al., 2001; Mauche et al., 2001; Doron & Behar, 2002; Xu et al., 2002; Gu, 2003; Paerels & Kahn, 2003; Werner et al., 2009; Pradhan & Nahar, 2011; Beiersdorfer et al., 2018; Gu et al., 2019, 2020; Grell et al., 2021).
@@
Despite decades of study, since early solar X-ray observations (Parkinson, 1973; Smith et al., 1985; Schmelz et al., 1992; Waljeski et al., 1994), discrepancies between observed and theoretical intensity ratios (Brown et al., 1998) have persisted. Early explanations invoking resonance scattering (McKenzie et al., 1980; Schmelz et al., 1992; Saba et al., 1999) found no confirmation in measurements with electron beam ion traps (EBITs) and tokamaks that also agreed with solar observations (Brown et al., 1998, 2001a, 2001b, 2006; Beiersdorfer et al., 2002; Gillaspy et al., 2011; Beiersdorfer et al., 2004, 2017; Shah et al., 2019). As optically thin laboratory plasmas are not subject to resonance scattering, indirect line formation mechanisms were suggested (Chen & Pradhan, 2002; Gu, 2003; Beiersdorfer et al., 2008, 2014, 2015; Shah et al., 2019; Grilo et al., 2021; Gu et al., 2020). An experiment with a free-electron laser aimed at directly determining the oscillator-strength ratio for lines 3C and 3D without uncertainties due to electron-impact excitation. Its unexpected results departing even more from theory were attributed to inaccuracies in calculated oscillator strengths (Bernitt et al., 2012), but soon after, unforeseen transient non-equilibrium effects and population transfer due to the ultra-brilliant peak photon flux explained them (Oreshkina et al., 2014, 2016; Loch et al., 2015; Wu & Gao, 2019). Our later measurements (Kühn et al., 2020) with synchrotron radiation avoided this non-linear systematic and improved the accuracy of the oscillator strength ratio while still disagreeing with theory. Finally, further increases in resolving power and signal-to-noise ratio found the cause of the persistent discrepancies in hitherto unresolvable line wings and diffraction effects, and brought the oscillator-strength ratio in line with state-of-the-art predictions (Kühn et al., 2022).
@@
In spite of these advances, many questions remain open for this essential ion and many other less-studied species. For instance, widely used wavelength references for Fe XVII from EBIT measurements using a crystal spectrometer with a resolving power of 500–700 have uncertainties of 1–3 mÅ ( 40–200 meV), Doppler-equivalent to 15–50 km/s, i.e. 60–180 parts-per-million (ppm), for , and double that for high- transitions (Beiersdorfer & Wargelin, 1994; Brown et al., 1998). This is only marginally adequate for analysis of high resolution diffraction grating spectra acquired with the Chandra High Energy Transmission Grating Spectrometer (HETGS), which can measure velocities of bright emission lines with 25 km/s systematic uncertainty (Ishibashi et al., 2006; Bozzo et al., 2023). These uncertainties in transition energies will also impair achievement of the science goals of other extant, upcoming, and proposed missions, including XMM-Newton (Jansen et al., 2001; den Herder et al., 2001), XRISM (Tashiro et al., 2018)), Athena (Pajot et al., 2018; Barret et al., 2016), LEM (Kraft et al., 2022), Arcus (Heilmann et al., 2022; Smith et al., 2022), and Lynx (Schwartz et al., 2019). Even though some of these missions feature spectrometers with full-width-at-half-maximum (FWHM) resolution far broader than the uncertainties of published rest-energy determinations, well-exposed spectra of bright objects with a high signal-to-noise ratio will allow centroid determination with uncertainties comparable to or smaller than these prior measurements. There is clearly a need for better determinations of the Fe XVII transition energies that will allow us to take full advantage of the resolving power of current and future missions, as well as improved and well-benchmarked theoretical methods that can provide energies for transitions which have not yet been measured with sufficient precision.
@@
We report new measurements of the rest energies of key Fe XVII transitions with an EBIT at the P04 beamline of the PETRA III synchrotron with uncertainties below 15 ppm, an improvement by a factor of 4-20 over the status quo. The accuracy of our results translates in velocity terms to 5 km/s, and fully unlocks the value of archived and forthcoming observations from XMM-Newton and Chandra, as well as of accurate velocimetry targeted by upcoming missions (Pajot et al., 2018; Barret et al., 2016; Kraft et al., 2022; Heilmann et al., 2022; Smith et al., 2022). We also test the large-scale configuration interaction (CI) approach and, therefore, our combination of the CI and coupled-cluster approaches (CI + all-order method), crucial for the development of high-precision clocks (Kozlov et al., 2018), and essential for understanding quantum electrodynamics (QED) effects in many-electron system. By applying the model potential approach (Tupitsyn et al., 2016) using the QEDMOD package (Shabaev et al., 2018), we incorporate QED effects into the effective Hamiltonian, basis-set orbitals, and one-electron matrix elements–a widely employed practice. The quality of the QED model potential is usually assessed against exact solutions for H-like ions, since the uncertainty in the electronic correlation in HCI with a few valence electrons is usually larger, or at the level of QED contributions, unless the ionization degree is rather high. Until this work, there were no estimates regarding the accuracy of QEDMOD approach for the majority of many-electron systems.
@@
Motivated by our highly-accurate experimental results, we carry out new CI computations, taking into account contributions from high states, increasing the number of relativistic configurations from 1.2 million in our previous work (Kühn et al., 2022; Cheung et al., 2021) to over 4 million, and investigating the convergence of the computations in both of these parameters. The results show a remarkable degree of numerical convergence across all energy levels, and agree with the measurements to a level of 1–33 meV (1–45 ppm) that is unprecedented for a complex ion such as Fe XVII. For the first time, uncertainties in the electronic correlations smaller than QED corrections allow us to test the accuracy of the QED contribution in a many-electron system.
@@ 2 Measurements and Data Analysis
PolarX-EBIT (Micke et al., 2018) was designed for the study of highly charged ions interacting with X-ray photons at synchrotrons and free-electron lasers (see Leutenegger et al., 2020; Togawa et al., 2020; Kühn et al., 2020, 2022; Steinbrügge et al., 2022; Stierhof et al., 2022). Its off-axis electron gun emits a nearly monoenergetic electron beam that is compressed to a diameter of less than 100 µm by a magnetic field of 870 mT generated by permanent magnets. Considering the overlap of the ion cloud and the electron beam leads to an effective electron density of . Iron pentacarbonyl (Fe(CO)5) molecules enter the trap region as tenuous beam through a two-stage differential pumping system. There, electron impact dissociation generates Fe atoms, and step-wise electron-impact ionization produces highly charged ions that remain radially trapped by the ensuing negative space-charge potential of the electron beam, and axially by biased cylindrical drift tubes. We chose operating conditions to ensure that Fe XVII ions mostly populate the trap.
@@
At the soft X-ray beamline P04, an APPLE II undulator (Viefhaus et al., 2013) produces circularly polarized photons, which are then sent through a monochromator equipped with a variable line-spacing grating of 1200 lines/mm mean groove density. Using an exit slit opening of 50 µm, the energy resolution E was set to a value of approximately E/E13,000 in the energy range of 700 to 1100 eV. A pair of plane-elliptical mirrors refocuses this beam onto the ion cloud. The photon beam energy is scanned over the Fe XVII transitions of interest and the corresponding calibration lines. Two silicon drift detectors (SDD) mounted at the top and on the side of the EBIT register fluorescence, with 100 eV FWHM resolution, following from resonant photoexcitation as well as electron-impact excitation.
@@
To calibrate the monochromator photon energy scale, we excite K-shell transitions in H-like and He-like oxygen, fluorine, and neon ions trapped in PolarX-EBIT. Their energies can be calculated with uncertainties well below 1 meV. We take values for the H-like transitions from Yerokhin & Shabaev (2015), and from Erickson (1977) for up to , and for He-like ions, we take energy values from Yerokhin & Surzhykov (2019) for transitions up to .
@@
The monochromator disperses the spectrum of the undulator cone on the exit slit by choice of incidence and diffraction angles of the grating, which is accomplished by appropriate rotations of both the grating and mirror, with the extra degree of freedom removed by requiring fulfillment of the constant fix-focus (CFF) condition (Follath, 2001). The absolute angles of both grating and mirror are recorded using angular encoders. To measure electronic transitions with narrow natural line widths ¡ 50 meV, angular increments as small as degrees have to be resolved, equivalent to 36 milliarcseconds or 175 nrad. The installed Heidenhain RON 905 angular encoders have 36 000 reference marks per turn, or one every 10-2 degrees (36 arcseconds). Such encoders interpolate angle changes a thousand times between each mark. A LED source is positioned on one side of the disk, while two photodiodes are positioned on the opposite side of the disk to record the light intensity modulated by slight rotations of the encoder disk111https://www.heidenhain.us/wp-content/uploads/2022/07/591109-24_Angle_Encoders_with_Integral_Bearing.pdf. These intensity variations are then stored in an empirical look-up table in the hardware. However, the process is highly sensitive to imperfections in the analog signals, which can lead to periodic subdivision errors (Krempaský et al., 2011; Follath & Balzer, 2010). Furthermore, within each monochromator, there are two encoders—one dedicated to the grating and the other to the focusing mirror. This doubles the uncertainty in the interpolation, impacting the determination of the diffraction angle. Consequently, this can cause the nominal monochromator energy to deviate from the actual photon beam energy. This problem was previously observed in our studies at P04 (Kühn et al., 2022; Togawa et al., 2023) and other beamlines (Follath & Balzer, 2010; Krempaský et al., 2011), and leads to periodic fluctuations in the nominal photon-energy scale, which in our case have peak-to-peak amplitudes of up to 50 meV below 900 eV and 70 meV above 900 eV.
@@
To correct for them while scanning the monochromator to excite resonant transitions, we direct the photon beam exiting the EBIT onto a gold target mounted on a high-resolution hemispherical electron-energy analyzer, ASPHERE (Rossnagel et al., 2001), as shown in Fig. 1. There, photoelectrons are emitted, and their kinetic energy measured. The kinetic energy of these photoelectrons is given by the difference between the photon energy and the binding energy of the electrons, along with any potential bias applied to the gold target. If the bias potential applied to the target is constant, any change in the photon energy will manifest itself as a change in the kinetic energy of the Au electrons. However, if we change the target bias to track changes in the nominal photon energy, the electron kinetic energy remains nominally constant and the photoelectron peak (Au ) can appear at a fixed position on the electron detector, see Fig. 1 (a). Thus, any deviation of the actual photon energy from the nominal photon energy set by the monochromator would result in a deviation of the kinetic energy of the photoelectrons. An example of such a deviation, reflecting the interpolation inaccuracies of the two angular encoders of the monochromator, is shown in Fig. 1 (b). We fit Au peaks (linewidths of about 700 meV FWHM) with Voigt profiles to find their centroids and determine their kinetic energies with uncertainties of a few meV at electron count rates of 10. By cooling the gold target to liquid nitrogen temperature, we further reduce the peak width to 450 meV, which further improves centroid determination. We then use this information to correct each step of the nominal monochromator energy scale. To avoid any assumptions in modeling these deviations shown in Fig. 1 (b), and because the addition of an arbitrary constant term to the energy scale will be removed when calibrating against known reference energies, we simply add the measured electron kinetic energy directly to the nominal monochromator energy scale rather than first subtracting a nominal kinetic energy offset. On this corrected monochromator energy scale (Fig. 1 (c)), we then determine the centroids of the calibration lines and associate them with the theoretical references mentioned above. By fitting a third-order polynomial to these data, we obtain the dispersion curve, thus the calibrated monochromator energy scale.
@@
For the Fe XVII measurements we set the EBIT to use a mA, 3500 eV electron beam, capable of directly exciting the lines studied here, and thus generating an undesired background. These parameters yielded a ratio of photoexcitation peak to electron-impact background between 2 and 3 throughout the experiment, indeed lower than the ratio of 45 achieved in our previous work (Kühn et al., 2022) by cyclically switching the electron-beam energy between ion breeding and probing energy after a long parameter optimization. This time, since switching tests showed a severe loss of Fe XVII ions, we decided to use a constant electron beam energy of 3500 eV, well above that of dielectronic recombination satellites. The present signal-to-noise ratio and resolving power of 13 000 were sufficient for our reported accuracy.
@@
The P04 monochromator was scanned over ranges covering (3G and 3F), (3C and 3D), (3A and 3B), and transitions (4C and 4D) of Fe XVII. Fluorescence was collected in the SDDs for 10–15 s at each monochromator step. The count rate for each transition is directly proportional to the respective oscillator strength, and we can see transitions with excitation rates about 4–80 times lower than that of the 3C transition. Scans of each line were therefore repeated as needed to obtain good statistics. This also yielded adequate statistics at each step for the Au photoelectron peak position determination needed for the nominal monochromator photon energy scale correction. To construct the spectrum for a single transition, as depicted in Fig. 2, all photons detected in the SDDs within a 50-eV region of interest centered around the expected energy are summed as a function of monochromator energy. A representative scan for each of these lines is shown in Fig. 2. The transition energies of the Fe XVII lines were determined using a maximum-likelihood fit of Voigt profiles added to the linear background term arising from electron-impact excitation using the cash statistic (Cash, 1979; Kaastra, 2017). The Voigt function is a convolution of Lorentzian and Gaussian functions. The Gaussian contributions to the linewidth arise from the limited resolution of the monochromator and the thermal motion of the ions (Hoesch et al., 2022). The Lorentzian width, as shown in Kühn et al. (2022), stems from the natural linewidth of the transition and a pseudo-Lorentz instrumental component due to X-ray diffraction at beamline components (Follath & Balzer, 2010). Given the possible energy-dependent contributions to linewidths from beamline components, we chose to leave all parameters of the Voigt profile unconstrained during our fitting procedure for determining the line centroids.
@@
Table 1 presents the results for eight Fe XVII lines and their associated uncertainties from errors in centroid determination of calibration lines, the dispersion fit 1- confidence band, and the centroid determination of each Fe XVII line, which is typically in the range of 1–3 meV. The total systematic uncertainties of the calibration are estimated to reach levels of 10–15 meV. As mentioned before, the angular encoder interpolation error induces oscillations of the nominal monochromator photon energy scale up to meV in the 650–1150 eV energy range. While accurate reference energies and the corrections from ASPHERE to the photon energy axis help mitigating these oscillations, 10–20% (7-14 meV) residual variations still remain in the corrected monochromator photon energy scale. A potential source of these could be the limited resolution of the Keithley 6517 voltage source biasing the gold target. Despite using a 7-digit calibrated voltmeter (Agilent 3458a), the voltage source operates in 5 mV steps within the 100 V range, limiting our electron kinetic energy measurements. Further systematics arise from the frequent switching of the voltage range of this bias supply needed to cover the monochromator range of 600 – 1150 eV, requiring separate calibration for each voltage range. Unfortunately, we could only calibrate the bias supply in a narrow 20-V range. Moreover, unmeasured fluctuations in the voltages applied to the inner and outer hemispheres of the electron spectrometer may have introduced additional systematic uncertainties. The dimensional stability of its electron-optics components is affected by thermal drifts caused by varying ambient conditions at ppm levels to which we are already sensitive. Note that we also fit our data with skewed Voigt profiles, allowing for a non-zero skewness, thus accounting for any line asymmetries that may exist due to monochromator imperfections (Perry-Sassmannshausen et al., 2021; Hoesch et al., 2022; Togawa et al., 2023). These tests resulted in changes to the line centroids of less than meV, which is negligible considering the total uncertainty of our measurements. We considered whether lines from contaminant ions of oxygen, fluorine, or neon originating from residual calibration gases could lead to systematic errors in any of our transition energy determinations. Because these lines have known transition energies, and because of the extremely high resolving power (13000) attained in our experiment, we ruled out any significant effect from such contamination. After conservatively considering all these sources, our present uncertainties are a factor of 4-20 smaller than those of previously reported experiments (Beiersdorfer & Wargelin, 1994; Brown et al., 1998).

### v3

# High-Precision Transition Energy Measurements of Neon-like Fe XVII Ions

@@ Abstract
We improve by a factor of 4–20 the energy accuracy of the strongest soft X-ray transitions of Fe XVII ions by resonantly exciting them in an electron beam ion trap with a monochromatic beam at the P04 beamline of the PETRA III synchrotron facility. By simultaneously tracking instantaneous photon-energy fluctuations with a high-resolution photoelectron spectrometer, we minimize systematic uncertainties down to 10–15 meV, or velocity equivalent 5 km s-1 in their rest energies, substantially improving our knowledge of this key astrophysical ion. Our large-scale configuration-interaction computations include more than four million relativistic configurations and agree with the experiment at a level without precedent for a 10-electron system. Thereby, theoretical uncertainties for interelectronic correlations become far smaller than those of quantum electrodynamics (QED) corrections. The present QED benchmark strengthens our trust in future calculations of many other complex atomic ions of interest to astrophysics, plasma physics, and for the development of optical clocks with highly charged ions.
@@ 1 Introduction
Over the past three decades, extensive research has focused on the soft X-ray emission from Ne-like iron (Fe XVII, Fe16+), particularly in hot astrophysical plasmas observed by Chandra and XMM-Newton (Behar et al., 2001; Brinkman et al., 2001). The dominant spectral transitions and of Fe XVII within the 700–850 eV range (14.5 to 17.5 Å) play a crucial role in deducing the plasma parameters across various sources. These parameters include the electron temperature, density, elemental abundance, gas motion, and photon scattering opacity (Parkinson, 1973; Smith et al., 1985; Schmelz et al., 1992; Waljeski et al., 1994; Phillips et al., 1996; Behar et al., 2001; Mauche et al., 2001; Doron & Behar, 2002; Xu et al., 2002; Gu, 2003; Paerels & Kahn, 2003; Werner et al., 2009; Pradhan & Nahar, 2011; Beiersdorfer et al., 2018; Gu et al., 2019, 2020; Grell et al., 2021).
@@
Despite decades of study, since early solar X-ray observations (Parkinson, 1973; Smith et al., 1985; Schmelz et al., 1992; Waljeski et al., 1994), discrepancies between observed and theoretical intensity ratios (Brown et al., 1998) have persisted. Early explanations invoking resonance scattering (McKenzie et al., 1980; Schmelz et al., 1992; Saba et al., 1999) found no confirmation in measurements with electron-beam ion traps (EBITs) and tokamaks that also agreed with solar observations (Brown et al., 1998, 2001a, 2001b, 2006; Beiersdorfer et al., 2002; Gillaspy et al., 2011; Beiersdorfer et al., 2004, 2017; Shah et al., 2019). As optically thin laboratory plasmas are not subject to resonance scattering, indirect line formation mechanisms were suggested (Chen & Pradhan, 2002; Gu, 2003; Beiersdorfer et al., 2008, 2014, 2015; Shah et al., 2019; Grilo et al., 2021; Gu et al., 2020). An experiment with a free-electron laser aimed at directly determining the oscillator-strength ratio for lines 3C and 3D without uncertainties due to electron-impact excitation. Its unexpected results departing even more from theory were attributed to inaccuracies in calculated oscillator strengths (Bernitt et al., 2012), but soon after, unforeseen transient nonequilibrium effects and population transfer due to the ultrabrilliant peak photon flux explained them (Oreshkina et al., 2014, 2016; Loch et al., 2015; Wu & Gao, 2019). Our later measurements (Kühn et al., 2020) with synchrotron radiation avoided this nonlinear systematic and improved the accuracy of the oscillator-strength ratio while still disagreeing with the theory. Finally, further increases in resolving power and signal-to-noise ratio found the cause of the persistent discrepancies in hitherto unresolvable line wings and diffraction effects and brought the oscillator-strength ratio in line with state-of-the-art predictions (Kühn et al., 2022).
@@
In spite of these advances, many questions remain open for this essential ion and many other less-studied species. For instance, widely used wavelength references for Fe XVII from EBIT measurements using a crystal spectrometer with a resolving power of 500–700 have uncertainties of 1–3 mÅ ( 40–200 meV), Doppler-equivalent to 15–50 km s-1, i.e. 60–180 parts per million (ppm), for , and double that for high- transitions (Beiersdorfer & Wargelin, 1994; Brown et al., 1998). This is only marginally adequate for analysis of high-resolution diffraction grating spectra acquired with the Chandra High Energy Transmission Grating Spectrometer (HETGS), which can measure velocities of bright emission lines with 25 km s-1 systematic uncertainty (Ishibashi et al., 2006; Bozzo et al., 2023). These uncertainties in transition energies will also impair the achievement of the science goals of other extant, upcoming, and proposed missions, including XMM-Newton (Jansen et al., 2001; den Herder et al., 2001), XRISM (Tashiro et al., 2018)), Athena (Pajot et al., 2018; Barret et al., 2016), LEM (Kraft et al., 2022), Arcus (Heilmann et al., 2022; Smith et al., 2022), and Lynx (Schwartz et al., 2019). Even though some of these missions feature spectrometers with full-width-at-half-maximum (FWHM) resolution far broader than the uncertainties of published rest-energy determinations, well-exposed spectra of bright objects with a high signal-to-noise ratio will allow centroid determination with uncertainties comparable to or smaller than these prior measurements. There is clearly a need for better determinations of the Fe XVII transition energies that will allow us to take full advantage of the resolving power of current and future missions, as well as improved and well-benchmarked theoretical methods that can provide energies for transitions that have not yet been measured with sufficient precision.
@@
We report new measurements of the rest energies of key Fe XVII transitions with an EBIT at the P04 beamline of the PETRA III synchrotron with uncertainties below 15 ppm, an improvement by a factor of 4–20 over the status quo. The accuracy of our results translates in velocity terms to 5 km s-1, and fully unlocks the value of archived and forthcoming observations from XMM-Newton and Chandra, as well as of accurate velocimetry targeted by upcoming missions (Pajot et al., 2018; Barret et al., 2016; Kraft et al., 2022; Heilmann et al., 2022; Smith et al., 2022). We also test the large-scale configuration interaction (CI) approach and, therefore, our combination of the CI and coupled-cluster approaches (CI + all-order method), which is crucial for the development of high-precision clocks (Kozlov et al., 2018), and essential for understanding the quantum electrodynamics (QED) effects in many-electron systems. By applying the model potential approach (Tupitsyn et al., 2016) using the QEDMOD package (Shabaev et al., 2018), we incorporate QED effects into the effective Hamiltonian, basis-set orbitals, and one-electron matrix elements—a widely employed practice. The quality of the QED model potential is usually assessed against exact solutions for H-like ions since the uncertainty in the electronic correlation in HCI with a few valence electrons is usually larger, or at the level of QED contributions unless the ionization degree is rather high. Until this work, there were no estimates regarding the accuracy of the QEDMOD approach for the majority of many-electron systems.
@@
Motivated by our highly accurate experimental results, we carry out new CI computations, taking the contributions from high states into account, increasing the number of relativistic configurations from 1.2 million in our previous work (Kühn et al., 2022; Cheung et al., 2021) to over 4 million, and investigating the convergence of the computations in both of these parameters. The results show a remarkable degree of numerical convergence across all energy levels and agree with the measurements to a level of 1–33 meV (1–45 ppm) that is unprecedented for a complex ion such as Fe XVII. For the first time, uncertainties in the electronic correlations smaller than QED corrections allow us to test the accuracy of the QED contribution in a many-electron system.
@@ 2 Measurements and Data Analysis
PolarX-EBIT (Micke et al., 2018) was designed for the study of highly charged ions interacting with X-ray photons at synchrotrons and free-electron lasers (see Leutenegger et al., 2020; Togawa et al., 2020; Kühn et al., 2020, 2022; Steinbrügge et al., 2022; Stierhof et al., 2022). Its off-axis electron gun emits a nearly monoenergetic electron beam that is compressed to a diameter of less than 100 µm by a magnetic field of 870 mT generated by permanent magnets. Considering the overlap of the ion cloud and the electron beam leads to an effective electron density of . Iron pentacarbonyl (Fe(CO)5) molecules enter the trap region as a tenuous beam through a two-stage differential pumping system. There, electron-impact dissociation generates Fe atoms, and step-wise electron-impact ionization produces highly charged ions that remain radially trapped by the ensuing negative space-charge potential of the electron beam, and axially by biased cylindrical drift tubes. We chose operating conditions to ensure that Fe XVII ions mostly populate the trap.
@@
At the soft X-ray beamline P04, an APPLE II undulator (Viefhaus et al., 2013) produces circularly polarized photons, which are then sent through a monochromator equipped with a variable line-spacing grating of 1200 lines mm-1 mean groove density. Using an exit slit opening of 50 µm, the energy resolution E was set to a value of approximately E/E13,000 in the energy range of 700–1100 eV. A pair of plane-elliptical mirrors refocus this beam onto the ion cloud. The photon beam energy is scanned over the Fe XVII transitions of interest and the corresponding calibration lines. Two silicon drift detectors (SDD) mounted at the top and on the side of the EBIT register fluorescence, with 100 eV FWHM resolution, following from resonant photoexcitation as well as electron-impact excitation.
@@
To calibrate the monochromator photon-energy scale, we excite K-shell transitions in H-like and He-like oxygen, fluorine, and neon ions trapped in PolarX-EBIT. Their energies can be calculated with uncertainties well below 1 meV. We take values for the H-like transitions from Yerokhin & Shabaev (2015), and from Erickson (1977) for up to , and for He-like ions, we take energy values from Yerokhin & Surzhykov (2019) for transitions up to .
@@
The monochromator disperses the spectrum of the undulator cone on the exit slit by choice of incidence and diffraction angles of the grating, which is accomplished by appropriate rotations of both the grating and mirror, with the extra degree of freedom removed by requiring fulfillment of the constant fix-focus (CFF) condition (Follath, 2001). The absolute angles of both grating and mirror are recorded using angular encoders. To measure electronic transitions with narrow natural line widths ¡ 50 meV, angular increments as small as degrees have to be resolved, equivalent to 36 milliarcseconds or 175 nrad. The installed Heidenhain RON 905 angular encoders have 36 000 reference marks per turn, or one every 10-2 degrees (36 arcseconds). Such encoders interpolate angle changes a thousand times between each mark. An LED source is positioned on one side of the disk, while two photodiodes are positioned on the opposite side of the disk to record the light intensity modulated by slight rotations of the encoder disk.111https://www.heidenhain.us/wp-content/uploads/2022/07/591109-24_Angle_Encoders_with_Integral_Bearing.pdf These intensity variations are then stored in an empirical lookup table in the hardware. However, the process is highly sensitive to imperfections in the analog signals, which can lead to periodic subdivision errors (Krempaský et al., 2011; Follath & Balzer, 2010). Furthermore, within each monochromator, there are two encoders–one dedicated to the grating and the other to the focusing mirror. This doubles the uncertainty in the interpolation, impacting the determination of the diffraction angle. Consequently, this can cause the nominal monochromator energy to deviate from the actual photon beam energy. This problem was previously observed in our studies at P04 (Kühn et al., 2022; Togawa et al., 2023) and other beamlines (Follath & Balzer, 2010; Krempaský et al., 2011), and leads to periodic fluctuations in the nominal photon-energy scale, which in our case have peak-to-peak amplitudes of up to 50 meV below 900 eV and 70 meV above 900 eV.
@@
To correct for them while scanning the monochromator to excite resonant transitions, we direct the photon beam exiting the EBIT onto a gold target mounted on a high-resolution hemispherical electron-energy analyzer, ASPHERE (Rossnagel et al., 2001), as shown in Fig. 1. There, photoelectrons are emitted, and their kinetic energy is measured. The kinetic energy of these photoelectrons is given by the difference between the photon energy and the binding energy of the electrons, along with any potential bias applied to the gold target. If the bias potential applied to the target is constant, any change in the photon energy will manifest itself as a change in the kinetic energy of the Au electrons. However, if we change the target bias to track changes in the nominal photon energy, the electron kinetic energy remains nominally constant and the photoelectron peak (Au ) can appear at a fixed position on the electron detector, see Fig. 1 (a). Thus, any deviation of the actual photon energy from the nominal photon energy set by the monochromator would result in a deviation of the kinetic energy of the photoelectrons. An example of such a deviation, reflecting the interpolation inaccuracies of the two angular encoders of the monochromator, is shown in Fig. 1 (b). We fit Au peaks (line widths of about 700 meV FWHM) with Voigt profiles to find their centroids and determine their kinetic energies with uncertainties of a few meV at electron count rates of 10. By cooling the gold target to liquid nitrogen temperature, we further reduce the peak width to 450 meV, which further improves centroid determination. We then use this information to correct each step of the nominal monochromator energy scale. To avoid any assumptions in modeling these deviations shown in Fig. 1 (b), and because the addition of an arbitrary constant term to the energy scale will be removed when calibrating against known reference energies, we simply add the measured electron kinetic energy directly to the nominal monochromator energy scale rather than first subtracting a nominal kinetic energy offset. On this corrected monochromator energy scale (Fig. 1 (c)), we then determine the centroids of the calibration lines and associate them with the theoretical references mentioned above. By fitting a third-order polynomial to these data, we obtain the dispersion curve and thus the calibrated monochromator energy scale.
@@
For the Fe XVII measurements, we set the EBIT to use a mA, 3500 eV electron beam, capable of directly exciting the lines studied here, and thus generating an undesired background. These parameters yielded a ratio of photoexcitation peak to electron-impact background between 2 and 3 throughout the experiment, indeed lower than the ratio of 45 achieved in our previous work (Kühn et al., 2022) by cyclically switching the electron-beam energy between ion breeding and probing energy after a long parameter optimization. This time, since switching tests showed a severe loss of Fe XVII ions, we decided to use a constant electron-beam energy of 3500 eV, well above that of dielectronic recombination satellites. The present signal-to-noise ratio and resolving power of 13,000 were sufficient for our reported accuracy.
@@
The P04 monochromator was scanned over ranges covering (3G and 3F), (3C and 3D), (3A and 3B), and transitions (4C and 4D) of Fe XVII. Fluorescence was collected in the SDDs for 10–15 s at each monochromator step. The count rate for each transition is directly proportional to the respective oscillator strength, and we can see transitions with excitation rates about 4–80 times lower than that of the 3C transition. Scans of each line were therefore repeated as needed to obtain good statistics. This also yielded adequate statistics at each step for the Au photoelectron peak position determination needed for the nominal monochromator photon-energy scale correction. To construct the spectrum for a single transition, as depicted in Fig. 2, all photons detected in the SDDs within a 50-eV region of interest centered around the expected energy are summed as a function of the monochromator energy. A representative scan for each of these lines is shown in Fig. 2. The transition energies of the Fe XVII lines were determined using a maximum-likelihood fit of Voigt profiles added to the linear background term arising from electron-impact excitation using the cash statistic (Cash, 1979; Kaastra, 2017). The Voigt function is a convolution of Lorentzian and Gaussian functions. The Gaussian contributions to the linewidth arise from the limited resolution of the monochromator and the thermal motion of the ions (Hoesch et al., 2022). The Lorentzian width, as shown in Kühn et al. (2022), stems from the natural linewidth of the transition and a pseudo-Lorentz instrumental component due to X-ray diffraction at beamline components (Follath & Balzer, 2010). Given the possible energy-dependent contributions to line widths from beamline components, we chose to leave all parameters of the Voigt profile unconstrained during our fitting procedure for determining the line centroids.
@@
Table 1 presents the results for eight Fe XVII lines and their associated uncertainties from errors in the centroid determination of calibration lines, the dispersion fit 1 confidence band, and the centroid determination of each Fe XVII line, which is typically in the range of 1–3 meV. The total systematic uncertainties of the calibration are estimated to reach levels of 10–15 meV. As mentioned before, the angular encoder interpolation error induces oscillations of the nominal monochromator photon energy scale up to meV in the 650–1150 eV energy range. While accurate reference energies and the corrections from ASPHERE to the photon-energy axis help mitigate these oscillations, 10–20% (7-14 meV) residual variations still remain in the corrected monochromator photon-energy scale. A potential source of these could be the limited resolution of the Keithley 6517 voltage source biasing the gold target. Despite using a seven-digit calibrated voltmeter (Agilent 3458a), the voltage source operates in 5 mV steps within the 100 V range, limiting our electron kinetic energy measurements. Further systematics arise from the frequent switching of the voltage range of this bias supply needed to cover the monochromator range of 600–1150 eV, requiring separate calibration for each voltage range. Unfortunately, we could only calibrate the bias supply in a narrow 20 V range. Moreover, unmeasured fluctuations in the voltages applied to the inner and outer hemispheres of the electron spectrometer may have introduced additional systematic uncertainties. The dimensional stability of its electron-optics components is affected by thermal drifts caused by varying ambient conditions at ppm levels to which we are already sensitive. Note that we also fit our data with skewed Voigt profiles, allowing for a nonzero skewness, thus accounting for any line asymmetries that may exist due to monochromator imperfections (Perry-Sassmannshausen et al., 2021; Hoesch et al., 2022; Togawa et al., 2023). These tests resulted in changes to the line centroids of less than meV, which is negligible considering the total uncertainty of our measurements. We considered whether lines from contaminant ions of oxygen, fluorine, or neon originating from residual calibration gases could lead to systematic errors in any of our transition-energy determinations. Because these lines have known transition energies, and because of the extremely high resolving power (13,000) attained in our experiment, we ruled out any significant effect from such contamination. After conservatively considering all these sources, our present uncertainties are a factor of 4–20 smaller than those of previously reported experiments (Beiersdorfer & Wargelin, 1994; Brown et al., 1998).